OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (10, –9); m = -2 a. y + 9 = -2(x – 10) b. y – 9 = -2(x - 10) c. y – 10 = -2(x + 9) d. y – 9 = -2(x + 10)

OpenStudy (anonymous):

point slope form is y - yo = m ( x - xo) plug in y-(-9) = -2 ( x-10) y+9 = -2(x-10) That is A